OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 6 2x − 4y = 10

OpenStudy (anonymous):

do you move the x to the other side first or do you just solve

OpenStudy (karatechopper):

Well have you heard of substitution or elimination?

OpenStudy (anonymous):

yes

OpenStudy (karatechopper):

Which process do you prefer?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

so the equation would look like this 2y=x-6

OpenStudy (karatechopper):

What variable do you want to use?

OpenStudy (anonymous):

x and y

OpenStudy (karatechopper):

Okay pick one. :)

OpenStudy (anonymous):

x-2y=6

OpenStudy (karatechopper):

Lol tyler. I'm asking you to pick a variable that we can solve for in order to substitute.

OpenStudy (anonymous):

oh lol x

OpenStudy (karatechopper):

ahah alrighty then.

OpenStudy (karatechopper):

x − 2y = 6 Solve that equation for x now. :)

OpenStudy (anonymous):

so you would substitute 0 for x and the equation (0)-2y=6

OpenStudy (karatechopper):

nooooooooooooo

OpenStudy (karatechopper):

where have you go that my friend

OpenStudy (karatechopper):

Solve for x :)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

x-2(0)=6

OpenStudy (anonymous):

x=6

OpenStudy (karatechopper):

.uhm. We need to isolate x x − 2y = 6 Add 2y to both sides in order to make x be on its own. Whats x= now?

OpenStudy (anonymous):

x=2y+6

OpenStudy (karatechopper):

YAY! :D

OpenStudy (karatechopper):

Okay now let's look at the second equation.

OpenStudy (karatechopper):

2x − 4y = 10

OpenStudy (anonymous):

ok

OpenStudy (karatechopper):

We know that x=2y+6 Now plug that value for x into this equation 2x − 4y = 10

OpenStudy (anonymous):

ok so 2(2y+6)-4y=10

OpenStudy (anonymous):

4y+12-4y=10

OpenStudy (anonymous):

8y+12=10 8y=-2

OpenStudy (karatechopper):

Wait a minute..

OpenStudy (karatechopper):

2(2y+6)-4y=10 4y+12-4y=10 4y-4y=0 12=10 Dang we did something wrong.

OpenStudy (anonymous):

yeah maybe we entered the wrong equation

OpenStudy (karatechopper):

uhm lets try solving for x in the bottom equation? 2x-4y=10

OpenStudy (karatechopper):

actually don't...it won't work again.

OpenStudy (karatechopper):

Are you okay if we try elimination?

OpenStudy (anonymous):

yeah sure

OpenStudy (karatechopper):

Yay. x − 2y = 6 2x − 4y = 10 Okay so let's try to eliminate x, we will multiply the top equation by -2, Could you do that for me please?

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

you multiply everthing by two or what

OpenStudy (karatechopper):

yes by -2

OpenStudy (anonymous):

ok so -2x-(-4)y=-20

OpenStudy (karatechopper):

what is 6 x -2?

OpenStudy (karatechopper):

We just wanted to multiply x-2y=6 You should have gotten (-2)x-(-2)(2)y=6(-2)

OpenStudy (anonymous):

-2x-(-4)y=-12

OpenStudy (karatechopper):

okay what is a negative times a negative

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

so its a positive 4y

OpenStudy (karatechopper):

yes

OpenStudy (karatechopper):

now rewrite the equation we have.

OpenStudy (anonymous):

-2x+4y=-12

OpenStudy (karatechopper):

Alright turn up.

OpenStudy (anonymous):

yay

OpenStudy (karatechopper):

So lets rewrite the whole thing. -2x+4y=-12 2x − 4y = 10 Now, solve it out. As you can see...both variables cancel out... Did you post the right question? Also, the answer because we cannot solve it is.... No solution.

OpenStudy (anonymous):

? Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x − 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x − y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (karatechopper):

Woah there amigo, what is this?

OpenStudy (anonymous):

the answers

OpenStudy (karatechopper):

What? Those look like dif questions, I am confused.

OpenStudy (anonymous):

ok look at this

OpenStudy (anonymous):

A system of equations is shown below: 6x − 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x − 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x − y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (anonymous):

thats the whole question

OpenStudy (karatechopper):

@TheSmartOne Will try to help you for a bit, I've got to finish something else. Hope I was some sort of help. Sorry.

OpenStudy (anonymous):

your fine you were help @karatechopper

TheSmartOne (thesmartone):

which question are we working on?

OpenStudy (anonymous):

the one up above

OpenStudy (anonymous):

but ill show you it just in case you didnt see it

OpenStudy (anonymous):

A system of equations is shown below: 6x − 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x − 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x − y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

TheSmartOne (thesmartone):

It's an interesting problem, hmmmm.

TheSmartOne (thesmartone):

can you add these 2 equations and tell me what you get: 6x − 2y = 3 5x + 3y = 4

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

so basically, what is 6x + 5x -2y + 3y = 3 + 4

OpenStudy (anonymous):

yeah sorry i was doing something else

OpenStudy (anonymous):

11x-y=7

TheSmartOne (thesmartone):

and which answer choice has 11x-y=7 ?

OpenStudy (anonymous):

the third and the fourth one

OpenStudy (anonymous):

C and D

TheSmartOne (thesmartone):

C has 11x \(\bf\color{red}{−}\) y = 7 D has 11x \(\bf \color{red}{+}\) y = 7

OpenStudy (anonymous):

so the answer is C

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

correct

OpenStudy (anonymous):

awesome thank you

TheSmartOne (thesmartone):

np :)